# Position: Many generalization measures for deep learning are fragile

### v2

@@ 4 Temporal behavior fragility
A plausible mechanism underlies this split. After interpolation under cross‑entropy, gradients can continue to increase logit scale along near‑flat directions. Adam tends to amplify this scale drift, while SGD+momentum tempers it, yielding opposite signs for the post‑interpolation slope. The lesson is that weight–norm–based measures—and margin‑normalized variants that co‑move with them—are not monotone indicators of optimization progress: an upward trend can be optimizer‑driven scale inflation; a flat or gently declining trend need not signal stagnation.
For practice, we recommend reporting $T_{\mathrm{int}}$ and the post‑interpolation slope $s\coloneqq\tfrac{d\log\mathrm{Measure}}{d\log t}\big|_{t>T_{\mathrm{int}}}$ alongside any temporal plot; overlay optimizers on the same axes; and compare snapshots at matched milestones (e.g., the first $100\%$‑accuracy crossing) or with early‑stopped checkpoints. A useful stress test is hysteresis: resume a checkpoint taken just after $T_{\mathrm{int}}$ with a different optimizer and check whether the measure’s drift changes sign while test error holds steady. These checks make temporal fragility visible and keep optimizer‑induced motion from being mistaken for learning. We refer the readers to Appendix B for experimental results on the temporal behaviors of more generalization measures.
@@ 5 Data-Complexity Fragility: Label Noise & Dataset Difficulty
There has been substantial recent interest in the study of scaling laws for deep neural networks (DNNs). These laws describe how generalization performance varies as a function of model size, training sample size, or computational cost (see, e.g., (Hestness et al., 2017; Kaplan et al., 2020; Spigler et al., 2019; Hoffmann et al., 2022; Valle-Pérez & Louis, 2020; Bahri et al., 2021; Nam et al., 2024; Goring et al., 2025) and references therein). Here, we observe that the ML-PACBayes bound does quite well at capturing the scaling behavior with the training sample size. However, it should be kept in mind that because this bound is evaluated with a Gaussian process (GP) method, it cannot capture feature-learning, which can play a crucial role in learning-curve scaling for certain datasets (Nam et al., 2024; Goring et al., 2025).
Figure 4: Fixed architecture, varying dataset.. The model is RESNET-50. Left: PACBAYES_ORIG with Adam. Middle: PACBAYES_ORIG with SGD (momentum). Changing from ADAM to SGD with momentum leads to a small increase in the overall scale of the PACBAYES_ORIG learning curves (solid lines). However, for both optimisers, these learning curves don’t distinguish MNIST, FashionMNIST, and CIFAR10, missing both the data‑complexity ordering and the changing slopes visible in the underlying learning curves (dashed lines). Right: ML‑PACBayes (marginal‑likelihood PAC–Bayes) is tight and preserves the correct dataset ordering and reproduces the trends in the slope; for technical reasons, this panel is computed on the binarized versions of the three datasets.
@@ 6 Scale-invariant network and exponential learning rate schedule
Normalization layers make many modern networks effectively scale‑invariant, and this symmetry underlies the “exponentially increasing learning‑rate” idea of Li & Arora (2019): in such models, training with fixed learning rate (LR), weight decay (WD), and momentum can be matched—in function space—by a run with an exponentially increasing LR and no WD (plus a small momentum correction across phases). Their analysis shows how LR and WD can be “folded into” an exponential schedule and documents successful experiments under this equivalence. We refer to this multiplicative control parameter $\alpha$ as the Exp++ factor.
We use the same symmetry but a different equivalence. Instead of removing WD, our theorem pairs an exponentially increasing LR with a time‑varying WD so that every iterate computes the same predictor as a fixed‑LR/fixed‑WD baseline. The mapping is explicit and non‑asymptotic: it gives closed‑form schedules for $\tilde{\eta}_{t}$ and $\tilde{\lambda}_{t}$, and an admissible range for the multiplicative factor $\alpha$ via the interval $\mathcal{I}$. Conceptually, this creates a single “Exp++” knob that can drive large changes in parameter norms while leaving the learned function essentially unchanged—exactly the benign intervention we later use to probe the fragility of magnitude‑sensitive measures. Empirical details and results are deferred to Appendix H.
Definition 6.1 (Scale invariant neural networks).
Theorem 6.2 (Equivalence of schedules in scale-invariant nets).
[…]
Define the interval endpoints
$\alpha_{L}$ $\coloneqq\frac{\gamma}{\,1-\lambda\eta_{0}+\gamma\,}$, (4)
and set
$\mathcal{I}\coloneqq(\alpha_{L},\ \alpha_{-}]\ \cup\ [\alpha_{+},\,1).$ (7)
If $\alpha\in\mathcal{I}$, then the schedules
$\tilde{\eta}_{t}$ $=\eta_{0}\,\alpha^{-2t-1},$ (8)
Remark 6.3 (Parameter range and nonnegativity).
Assume
$\lambda\eta_{0}$ $\leq(1-\sqrt{\gamma})^{2},$ (10)
Consequences. The proof of Theorem 6.2 is in Appendix C. In the scale‑invariant setting, this equivalence lets us pump parameter norms without changing $f(\theta)$, a lever we use to stress‑test generalization measures. Applying this dial, we find that even the measure that looked most stable in our earlier probes—the standard parameter‑space PAC–Bayes proxy PACBAYES_origfails catastrophically here: as the Exp++ factor grows, the proxy inflates by orders of magnitude while test error is essentially unchanged; see Appendix H.
@@ 7 Quantitative fragility: a scale‑free instability score
The heuristic stress tests in Sections 35 show that several popular post‑mortem measures are fragile. Here we provide a complementary quantitative lens: scale‑free statistics that capture how much a measure fluctuates among runs with (nearly) the same test error. These quantitative metrics of fragility can facilitate larger-scale comparisons between measures.
Setup.
Definition 7.1 (Conditional Measure Spread (CMS)).
The CMS of $C$ in group $g$ at tolerance $\delta$ is
Randomness adjustment.
Define the corresponding spreads
Definition 7.2 (eCMS (Excess CMS)).
The seed‑adjusted excess at matched error is
Aggregation across groups.
How to read the scores.
Properties and protocol.
Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$). Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (12)–(15) and the fraction of groups covered. Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters. Full experimental details, formal algorithms and tabulated results can be found in Appendix E.
8 Conclusion and discussion
We take a pragmatic view of generalization bounds: in modern deep learning they function best as measures to be judged by how they behave across data and training regimes rather than by worst‑case tightness. Read this way, fragility becomes a key property to audit. With the task fixed and accuracy essentially steady, small and reasonable pipeline tweaks—changing the learning rate by a factor of ten, swapping Adam for momentum SGD, toggling early stopping—can bend surrogate curves, invert trends, or inflate their scale. When the task complexity changes, via label corruption or by moving from MNIST to FashionMNIST to CIFAR‑10, several surrogates fail to preserve the expected ordering or slope. Our scale‑invariance construction makes this failure mode explicit: one can inflate parameter magnitudes without changing the predictor, and magnitude‑sensitive post‑mortems then balloon while test error does not. This gap underscores that properties of a single parameter vector need not reflect properties of the learned function.
The comparison between post‑mortem diagnostics and the marginal‑likelihood PAC–Bayes route in function space sharpens the lesson. This predictor can capture key trends with data complexity, including the slopes of learning curves. Given that the interaction between data and architecture is one of the key components of DNN performance that needs explaining, the success of this bound, and its simple interpretation in terms of inductive bias at initialisation, provides a concrete foil against which parameter‑space surrogates can be refined.
Looking forward, several concrete directions could strengthen this position:
  • Design reparameterization‑invariant, data‑aware post‑training diagnostics that borrow the Occam flavour of marginal likelihood while remaining practical at scale;
  • build a public fragility benchmark and automated audit harness covering optimizer/schedule swaps, scale‑symmetry probes, and data‑complexity sweeps.
  • Develop approximate function‑space surrogates (finite‑width corrections, amortized evidence estimators, ensembles as posteriors over functions) to make invariance‑friendly predictors usable in routine training.
  • Tie diagnostics to training dynamics by discouraging scale‑only drift after interpolation or by incorporating function‑space objectives during training.
  • Extend evaluation beyond vision classification to sequence models, generative modelling, reinforcement learning, and explicit distribution shift, testing whether the same invariance principles hold.
@@ 9 Alternative Views
  • Optimizer sensitivity is signal, not noise. Because implicit bias drives which predictor is learned, a useful measure should move with the training pipeline, not be invariant to it.
  • Post‑mortem measures can be stabilized. With the right reparameterization‑aware definitions (e.g., normalized sharpness, margin distributions, distance from initialization), much of the apparent fragility disappears.
  • Function‑space GP baselines are mis‑specified for modern, finite‑width practice. GP priors, absence of data augmentation in the prior, and finite‑width effects limit how decisively marginal‑likelihood predictors should be used as a reference.
Optimizer sensitivity as a feature.
It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalization measures. This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (Soudry et al., 2018; Smith et al., 2017; Simsekli et al., 2019). From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
@@
Stabilizing post‑mortem diagnostics.
A second objection holds that many post‑training measures already address scale and parameterization issues. Normalized or reparameterization‑aware sharpness correlates more reliably with generalization than raw curvature (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022); modeling the distribution of margins rather than the minimum improves predictiveness across runs (Jiang et al., 2018); reference‑aware surrogates such as distance from initialization or movement from pretraining reduce width and scale artifacts and come with supporting analyses (Li et al., 2018; Neyshabur et al., 2017a; Zhou et al., 2021). On this view, fragility largely reflects naive implementations, not intrinsic flaws.
@@
A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix F) overstates its practical authority. Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination. While GPs are known to closely match DNN performance on the datasets we test in this paper (Lee et al., 2020b), it is hard to see how one could untangle the effects of hyperparameters or optimiser choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address. More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified. In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys. Therefore, the GP route’s stability could stem from omitting factors that matter in practice (Lee et al., 2020b; Valle-Pérez & Louis, 2020).
@@
Response. We agree that one of the most important questions that needs explaining in DNNs is how, when, and why finite‑width networks outperform their infinite-width GP limits, in particular for settings where feature-learning is important (see e.g. (Goring et al., 2025) and references therein). We use the GP based marginal‑likelihood bound as a calibration tool, not an oracle: it naturally encodes desirable invariances (insensitivity to parameter scale and optimizer path once the dataset is fixed) and for the cases we study, consistently tracks data difficulty and learning‑curve scaling across families. Those properties make it a valuable foil for stress‑testing post‑mortem bounds. Where the GP prior is clearly misspecified (e.g., heavy augmentation or domains where feature-leaning is key), our recommendation is empirical: rerun the same fragility audit. If the GP predictor ceases to track error while a reparameterization‑invariant post‑mortem does, that is evidence for the importance of a post‑mortem analysis in that regime.
Synthesis.
These counterarguments motivate a middle path. Rather than elevating any single diagnostic, we advocate (i) reporting multiple invariance‑aware post‑mortem measures, (ii) auditing them with our stressors (learning‑curve shape, temporal drift after interpolation, and data‑complexity response), and (iii) anchoring interpretation with at least one function–space reference when feasible. This recognizes optimizer‑dependent mechanisms while discouraging conclusions drawn from features (e.g., raw scale) that can be changed without affecting predictions.
Impact Statement
This is a position paper about evaluation practice. The main impact is methodological: a simple audit protocol and reporting guidelines that encourage robustness checks for generalization measures. Potential positive impacts include more reliable model selection and clearer communication of limitations; potential negative impacts are minimal, as we do not propose deployable systems or datasets.
@@ Appendix A Training-hyperparameter fragility for all measures
In this section we give concise definitions of the measures we study (aligning notation with prior large-scale evaluations) and provide additional experimental evidence beyond the main text (Dziugaite et al., 2020b; Jiang et al., 2019b). We then present figure-backed comparisons where small learning-rate or optimizer tweaks trigger qualitatively different curve shapes for the surrogate—plateaus, rebounds, late spikes, and crossings—while the accompanying accuracy curves remain comparatively calm.

### v3

@@ 4 Temporal behavior fragility
A plausible mechanism underlies this split. After interpolation under cross‑entropy, gradients can continue to increase logit scale along near‑flat directions. Adam tends to amplify this scale drift, while SGD+momentum tempers it, yielding opposite signs for the post‑interpolation slope. The lesson is that weight-norm-based measures—and margin-normalized variants that co-move with them—are not monotone indicators of optimization progress: an upward trend can be optimizer‑driven scale inflation; a flat or gently declining trend need not signal stagnation. We refer the readers to Appendix B for experimental results on the temporal behaviors of more generalization measures.
@@ 5 Data-Complexity Fragility: Label Noise & Dataset Difficulty
There has been substantial recent interest in the study of scaling laws for deep neural networks (DNNs). These laws describe how generalization performance varies as a function of model size, training sample size, or computational cost (see, e.g., (hestness2017deep; kaplan2020scaling; spigler2019asymptotic; hoffmann2022training; valleprez2020generalization; bahri2024explaining; nam2024visualising; goring2025feature) and references therein). Here, we observe that the ML-PACBayes bound does quite well at capturing the scaling behavior with the training sample size. However, it should be kept in mind that because this bound is evaluated with a Gaussian process (GP) method, it cannot capture feature-learning, which can play a crucial role in learning-curve scaling for certain datasets (nam2024visualising; goring2025feature).
Figure 4: Fixed architecture, varying dataset.. The model is RESNET-50. Left: PACBAYES_ORIG with Adam. Middle: PACBAYES_ORIG with SGD (momentum). Changing from ADAM to SGD with momentum leads to a small increase in the overall scale of the PACBAYES_ORIG learning curves (solid lines). However, for both optimizers, these learning curves don’t distinguish MNIST, FashionMNIST, and CIFAR10, missing both the data‑complexity ordering and the changing slopes visible in the underlying learning curves (dashed lines). Right: ML‑PACBayes (marginal‑likelihood PAC–Bayes) is tight and preserves the correct dataset ordering and reproduces the trends in the slope; for technical reasons, this panel is computed on the binarized versions of the three datasets.
@@
6 Fragility of tightness: the case of compression bounds
Recent work by lotfi2022pacb proposes PAC-Bayes compression bounds that achieve state-of-the-art tightness. They argue that this tightness allows them to “explain” generalization. While this approach is theoretically stimulating, our audit suggests that numerical tightness does not necessarily imply a correct causal explanation of the generalization phenomenon. We identify two specific fragilities.
(a) CIFAR-10 (ID 3500) Bounds
(b) MNIST (ID 1000) Bounds
(c) KL Divergence (CIFAR vs MNIST)
Figure 5: Decomposing PAC-Bayes Compression Bounds. Left & Middle: Learning curves for bounds and error rates. The “quantized training error” (dashed brown) is large and tracks the test error. Right: The Raw KL divergence (complexity term) remains essentially flat for both datasets across sample sizes ($n$). Additional results for higher intrinsic dimensions are provided in Appendix I.
Bias dominance over complexity. In standard learning theory, bounds typically trade off a small training error against a complexity penalty. However, as shown in the left and middle panels of Figure 5, the “quantized training error” (dashed brown line)—which serves as the empirical risk term in the bound—is notably large. Crucially, as the sample size $n$ increases, this quantized training error does not vanish but instead grows to track the quantized test error. This suggests that the final bound’s tightness is driven primarily by this large bias term (the quantization error) rather than by the KL-divergence complexity term measuring the neural network’s capacity. In practice, modern DNNs often achieve near-zero training error; a bound that relies on a large proxy training error to achieve tightness is arguably measuring the quality of the compression scheme rather than the generalization mechanism of the original network.
Insensitivity of the complexity term. A generalization measure should ideally reflect how the difficulty of the learning problem changes with data availability. However, the right panel of Figure 5 demonstrates that the KL divergence—the core measure of complexity in the PAC-Bayes formulation—remains almost perfectly flat with respect to training set size $n$. This flatness persists across different datasets (CIFAR-10 vs. MNIST) and intrinsic dimensions (see Appendix I for results with $\text{ID}=10000$). This implies that the complexity term fails to capture the scaling of learning difficulty, rendering the bound insensitive to the actual information gain from increasing data.
@@ 7 Scale-invariant network and exponential learning rate schedule
Normalization layers render many modern networks effectively scale-invariant, motivating the exponentially increasing learning-rate scheme of li2019an. In Appendix C, we prove a non-asymptotic equivalence: training with fixed learning rate (LR) and weight decay (WD) is exactly matched in function space by an exponentially increasing LR with time-varying WD.
We refer to this multiplicative control parameter $\alpha$ as the Exp++ factor. This theoretical result provides a precise “invariance lever”: it allows us to arbitrarily inflate parameter norms while leaving the learned function $f(\theta)$—and thus the test error—mathematically identical at every step. This mechanism serves as a crucial control for our audit: if a generalization measure is sensitive to parameter magnitude (like path norm or standard PAC-Bayes), it will explode under this transformation, even though generalization performance remains identical.
The formal definitions, the statement of Theorem C.2, and its proof are provided in Appendix C. Empirical results utilizing this lever are detailed in Appendix H.
@@ 8 Quantitative fragility: a scale‑free instability score
The heuristic stress tests in Sections 35 show that several popular post‑mortem measures are fragile. To facilitate large-scale comparisons, we developed a complementary quantitative framework that assigns a specific instability score to a measure. We introduce two statistics: Conditional Measure Spread (CMS), which captures the fluctuation of a measure across runs with nearly identical test error, and Excess CMS (eCMS), which isolates the component of that fluctuation driven specifically by hyperparameter changes (net of random seed variance).
Summary of results. As detailed in Appendix E, we computed these scores for a wide range of measures. We found that parameter-count proxies are stable (eCMS $\approx 0$), while path norms, spectral distances, and standard PAC-Bayes surrogates exhibit high instability (large eCMS), quantitatively confirming the qualitative fragility observed in our stress tests. See Appendix E for the formal definitions (Definitions E.1 and E.2) and the full tabulated results.
@@ 9 Alternative views
Optimizer sensitivity as a feature. It is widely thought that optimizer, schedule, and batch size are first‑order determinants of generalization that should be reflected in generalization measures. This view points to well‑documented implicit‑bias phenomena—e.g., margin growth under separable losses and optimizer‑dependent convergence paths—that influence which classifier is ultimately selected (soudry2018the; smith2017donb; simsekli2019tail). From this perspective, the very invariances we prize risk washing out real, practically actionable differences between training recipes.
@@
Stabilizing post‑mortem diagnostics. A second objection holds that many post‑training measures already address scale and parameterization issues. Normalized or reparameterization‑aware sharpness correlates more reliably with generalization than raw curvature (dinh2017sharpb; tsuzuku2020normalized; kim2022scale; jang2022reparametrization); modeling the distribution of margins rather than the minimum improves predictiveness across runs (jiang2018predicting); reference‑aware surrogates such as distance from initialization or movement from pretraining reduce width and scale artifacts and come with supporting analyses (li2018on; neyshabur2017exploring; zhou2021understanding). On this view, fragility largely reflects naive implementations, not intrinsic flaws.
@@
Limits of function–space GP references. A third objection is that using a marginal‑likelihood PAC–Bayes predictor as a baseline (Appendix F) overstates its practical authority. Since a GP is used to approximate the key marginal-likelihood term in the bound, the ML-PACBayes predictor could be seen as a measure of the performance of the underlying GP for any given architecture-data combination. While GPs are known to closely match DNN performance on the datasets we test in this paper (lee2019wide), it is hard to see how one could untangle the effects of hyperparameters or optimizer choice in this approach, in other words, the kinds of questions that post-mortem bounds are trying to address. More generally, modern deep-learning models are finite‑width, heavily augmented, and often trained far from the GP regime; priors that ignore augmentation, architectural quirks, or fine‑tuned tokenization may be badly mis‑specified. In addition, the bound controls the error of a stochastic Gibbs classifier drawn from a posterior over functions, not necessarily the deterministic network one actually deploys. Therefore, the GP route’s stability could stem from omitting factors that matter in practice (lee2019wide; valleprez2020generalization).
@@
Response. We agree that one of the most important questions that needs explaining in DNNs is how, when, and why finite‑width networks outperform their infinite-width GP limits, in particular for settings where feature-learning is important (see e.g. (goring2025feature) and references therein). We use the GP based marginal‑likelihood bound as a calibration tool, not an oracle: it naturally encodes desirable invariances (insensitivity to parameter scale and optimizer path once the dataset is fixed) and for the cases we study, consistently tracks data difficulty and learning‑curve scaling across families. Those properties make it a valuable foil for stress‑testing post‑mortem bounds. Where the GP prior is clearly misspecified (e.g., heavy augmentation or domains where feature-learning is key), our recommendation is empirical: rerun the same fragility audit. If the GP predictor ceases to track error while a reparameterization‑invariant post‑mortem does, that is evidence for the importance of a post‑mortem analysis in that regime.
10 Call to action: a better standard for evaluation
Our position is not merely that current measures are fragile, but that the community must adopt new standards to diagnose and mitigate this fragility. We propose three concrete actions for authors, reviewers, and benchmark designers.
1. Mandate the Fragility Audit. Just as error bars are expected for experimental results, a “stability check” should accompany any proposed generalization measure. Authors introducing a new bound or proxy should report its behavior under the Minimum Fragility Stress-Test:
  • The Hyperparameter Sweep: Plot the measure against learning rate and optimizer changes. If the measure shifts by orders of magnitude while test error remains constant (as seen in Fig. 1), this sensitivity must be reported.
  • The Data-Complexity Check: Demonstrate that the measure correctly ranks datasets by difficulty (e.g., MNIST vs. CIFAR-10) or tracks label noise (Fig. 3).
  • The Temporal Check: Report the measure’s post-interpolation slope. A valid generalization measure should not drift indefinitely after the training error has hit zero.
2. Prioritize Invariance in Design. The community should pivot from deriving tighter worst-case bounds to designing invariant measures. We recommend developing post-mortem diagnostics that are explicitly invariant to layerwise reparameterizations and scale transformations (such as the Exp++ mechanism in §7). Measures based on raw parameter magnitudes should be avoided in favor of relative or normalized metrics, unless scale dependence is explicitly justified.
3. Develop Function-Space Surrogates. Our results with the ML-PACBayes bound suggest that function-space quantities naturally avoid parameter-space fragility. A major open challenge is to develop computable approximations of function-space complexity—such as approximate marginal likelihood or function-space entropy—that scale to modern deep networks. We call for a research program dedicated to estimating these quantities efficiently, potentially bridging the gap between the robustness of GP limits and the feature-learning capabilities of finite-width networks.
Conclusion. In summary, we have argued that generalization measures in deep learning must be judged not only by their tightness but also by their robustness. Through systematic auditing, we demonstrated that many popular post-mortem bounds—including path norms, spectral measures, and standard PAC-Bayes surrogates—are fragile: they can be made to vary wildly, trend in reverse, or collapse entirely under benign training changes that leave the underlying network’s performance effectively unchanged. More subtle, but important, fragilities include failing to capture basic generalisation trends with data size or data complexity.
This fragility is likely more pronounced for measures that probe parameter space rather than function space. By contrast, a simple function-based marginal-likelihood PAC-Bayes bound, which captures the inductive bias of the architecture, tracks data complexity and learning-curve scaling, though it cannot distinguish optimizer effects. Its performance provides a natural benchmark for future, more sophisticated measures that also capture post-training effects.
We contend that the current practice of evaluating post-training diagnostics—often limited to mild variations in training pipelines—is insufficient to expose true fragility. By adopting the more rigorous audits and invariance principles proposed in this work, the field can move toward measures that genuinely reflect and, more importantly, help us understand the mechanisms of generalization.
@@ Appendix A Training-hyperparameter fragility for all measures
In this section we give concise definitions of the measures we study (aligning notation with prior large-scale evaluations) and provide additional experimental evidence beyond the main text (dziugaite2020in; jiang2019fantastic). We then present figure-backed comparisons where small learning-rate or optimizer tweaks trigger qualitatively different curve shapes for the surrogate—plateaus, rebounds, late spikes, and crossings—while the accompanying accuracy curves remain comparatively calm.